# Quantum coherence and the invisible Universe: Subradiance as a dark matter mechanism

### v3

@@ 2.1 Entangled quantum states and coherent behavior
Following Dicke [1], the kets |$\Ket{r,m}$ are defined with the “cooperative” number $r$ and “inversion” number $m$, which are analogous to spin-1/2 quantum numbers, with $r\geq 0$ and $\left|m\right|\leq r$ (|$\Ket{r,m}_{\theta}=\Ket{r,m}$ when $\cos\theta=0$). Note that $k=2\pi/\lambda$ with $\lambda$ the radiation wavelength.
@@
The two-atom system will, on average, emit a photon half of the time but at twice the single-atom rate.
Another way to look at this is by considering the energy trapping efficiency defined as
@@
The calculations can be extended to a larger number of atoms. The internal states for an arbitrary number of atoms $n$ is shown in Figure 1 along with the expected transition rates in the small sample (i.e., $kz_{0}\rightarrow 0$) or coherent limits (taken from Dicke’s original paper [1]). It is generally observed that, starting from a given $m$ value, transitions are allowed only between states sharing the same cooperation number $r$, while rates systematically decrease as one moves rightward to lower $r$ values.
@@ 2.3 Superabsorption, coherence and triggered subradiance
We now study the propagation of an incident radiation field on a slab composed of an atomic gas, similar to what is often done in elementary radiative transfer analysis in astrophysics [23]. This will allow us to compare the outcome of our analysis with well-known results. The MBE are perfectly suited for this problem, being one-dimensional in nature. We will also investigate effects the incident field will have on the radiation emanating from within the gas itself. For a magnetic dipole transition, such as the atomic hydrogen 21 cm line, the MBE allow us to follow the temporal evolution of three quantities as a function of position $z$ and the retarded time $\tau=t-z/c$: the differential density $n_{-}=n_{b}-n_{a}$, with $n_{a}$ and $n_{b}$ are the number densities of atoms in the lower and upper levels [^2] From now on we use lower $n$’s for densities and capital $N$’s for numbers of particles., respectively, the magnetization $M^{+}$ and the magnetic field $B^{+}$. Our MBE include relaxation and dephasing time-scales $T_{1}$ and $T_{2}$, as well as a thermal equilibrium condition enforced through a corresponding differential density $n_{-}^{\mathrm{eq}}=n_{b}^{\mathrm{eq}}-n_{a}^{\mathrm{eq}}<0$ set by the Boltzmann equation for the population of energy levels at the temperature of the gas $T_{\mathrm{kin}}$.
@@ 2.3.1 Response to incident radiation
$\left.+\frac{\alpha}{2}e^{-\tau/T_{2}}\left[e^{-\alpha z/2}\star J_{0}\left(2\sqrt{\frac{z\tau}{LT_{R}}}\right)\right]\right\},$ (12)
@@
As we will soon be interested in the duration of the transient response of the system when $T_{2}\ll T_{R}$, we note that in this limit the evolution time-scale of the system is (see Appendix C.2)
$T_{\mathrm{tr}}\approx T_{2}\left(1+\sqrt{\frac{zT_{2}}{LT_{R}}}\right)^{-1}.$ (15)
The transient signal thus dies out on a time-scale of the order of $T_{\mathrm{tr}}$.
@@
$\alpha\approx\alpha_{\mathrm{B}}\left(1+\sqrt{\frac{zT_{2}}{LT_{R}}}\right).$ (22)
for $T_{2}\ll T_{R}$.
We uncover the important result that Beer’s law holds strictly only under conditions where $T_{2}\rightarrow 0$, i.e., in the limit of complete decoherence/dephasing. In the opposite limit when $T_{2}\gg T_{R}$, we have $\delta\omega\sim T_{R}^{-1}$, $\alpha=\alpha_{\mathrm{B}}T_{2}/T_{R}$ and find, just as importantly, that the density of the gas determined with Beer’s law can be overestimated when coherence is present.
@@
Figure 6: Left: Temporal evolutions of the radiation intensity (top) and magnetic induction field (bottom) of the 21 cm line for an atomic hydrogen gas at a temperature of $10$ K and length $L\simeq 1.5$ pc. We set $T_{1}=T_{2}\simeq 16\,T_{R}$ and the intensity is normalized to the saturation intensity $I_{\mathrm{sat}}$ given in equation (18). In the bottom panel, the incident magnetic field’s amplitude is shown in cyan (using the vertical scale on the right), while a broken curve in red corresponding to equation (12) is plotted on top of that obtained from numerical computations (in black) using the MBE. Right: Temporal evolution of the density $n_{-}$ in black with the pump $\Lambda_{n}=n_{-}^{\mathrm{eq}}/T_{1}$ in cyan (using the vertical scale on the right; top) and the magnetization (bottom). The density is limited to the spectral extent of the signal (i.e., $T_{R}^{-1}$) and corresponds to a total density of $1\,\mathrm{cm^{-3}}$ for a gas at a temperature of $10$ K. The mean distance between neighboring cooperating atoms is $\Delta r\approx 670\lambda$ and $T_{R}=1.5\times 10^{8}\,\mathrm{s}\simeq 4.3\times 10^{-7}/\Gamma$.
@@
The left panel of Figure 6 shows the temporal evolutions of the radiation intensity (normalized to $I_{\mathrm{sat}}$; top) and the magnetic field (bottom). A broken curve in red corresponding to equation (12) is plotted on top of that obtained from numerical computations using the MBE for the magnetic field (in black); there is perfect agreement between the two. The (constant) density $n_{-}$ (in black) and pump $\Lambda_{n}$ (in cyan; using the vertical scale on the right; top) and the magnetization (in units of Bohr magneton $\mu_{\mathrm{B}}$ per cubic centimeter; bottom) are displayed in the right panel. We note that the mean distance between neighboring cooperating atoms is $\Delta r\approx 670\lambda$ and $T_{2}\simeq 16T_{R}$, which implies that coherence can be sustained in the system under these conditions.
It is interesting to note that at $\tau=0$ the incident magnetic field is transmitted unattenuated at $z=L$ (or any other position) no matter how large $L$ is. Thereafter, superabsorption defines the transient response to greatly reduce the intensity on a time-scale $\approx T_{R}$ and in a oscillatory manner reminiscing of superradiance systems. According to our previous discussion, the steady-state regime settles afterwards to an intensity diminished by a factor $e^{-\alpha L}$, where $\alpha L=2T_{2}/T_{R}\simeq 32$ from equation (13). Such signal would therefore be detected as a deep and narrow absorption line ($\approx 0.5\,\mathrm{km\,s^{-1}}$ in width) against the background radiation.
@@ 2.3.2 Response to initial fluctuations
More precisely, at $\tau=0$ the system’s intensity is at a level similar to what is expected from a non-coherent gas (i.e., $I=0.25\,I_{\mathrm{nc}}$), but it only radiates for a duration $\approx T_{R}$. To that effect, we should emphasize that $\Gamma T_{R}\simeq 4.3\times 10^{-7}$. In comparison, for the 21 cm line the non-coherent transient intensity for non-cooperating atoms would last on the order of $\Gamma^{-1}\sim 10^{7}$ yr, whereas $T_{R}\sim 5$ yr. This implies that almost the totality of the available energy is locked within the atomic population, which is almost entirely in dark states since the fast states are quickly depleted during the transient regime.
@@ 3 Discussion
In our analysis of the 21 cm line in Sec. 2.3, we purposely chose a gas density that is in line with measurements in dark matter halos. For the matter contained in these environments, the main source of anisotropy comes in the form of the radiation field emanating from the host galaxy. As then mentioned, the incidence of this field will effectively select one mode for the propagation of radiation within the halo. It is expected that superabsorption of the incident field will elicit coherence in subsequent transitions in the gas through angular correlation among photons. As a consequence, extreme subradiance, leading to a complete darkening of the gas, takes place within a few years even at a mean separation $\Delta r\approx 670\lambda$ between closest neighbors. As far as emission properties are concerned, this atomic hydrogen gas is invisible and effectively behaves like dark matter. This darkening effect may therefore possibly account for the dark matter known to exist in these environments.
@@
For our analysis of the 21 cm line for a gas at a density $n_{-}^{\mathrm{eq}}=1\,\mathrm{cm^{-3}}$ and temperature $T_{\mathrm{kin}}=10$ K these considerations have no bearing on the outcome since all other lines (e.g., Lyman-$\alpha$) will remain essentially unexcited. The gas as a whole will thus appear dark in the steady-state on the account of the 21 cm line alone. While we only studied this specific example, similar conclusions may be reached when considering other environments with differing gas densities and temperatures, or other atomic/molecular species. It also follows that this darkening effect, which also relies on the presence of anisotropic radiation, should be closely correlated with the presence of “luminous” matter, as is the case for dark matter [27].
@@
We are then led to a picture where our “dark” hydrogen gas (in emission) would also appear more opaque (or more optically thick) in its absorption lines against the background source responsible for the incident radiation. However, it would otherwise be completely transparent to radiation at wavelengths alien to the hydrogen atom, as expected from dark matter. For a cold gas as the one considered in our analysis, these lines would appear as narrow absorption features, such as those commonly detected in observations of the interstellar medium [28], the Galactic halo [29], the Galactic disk [30] and external galaxies [31].
Another important property of our darkened atomic hydrogen gas is discussed in Appendix D. There, we calculate the cross-section for high-speed collisions between two two-atom systems (i.e., when the kinetic energy is far greater than the transition excitation energy $\hbar\omega_{0}$), both being entangled as discussed in Sec. 2.1. Interestingly, it is found that in such circumstances the collision cross-section will likely vanish because of destructive interference due to the entangled nature of the states the systems find themselves in after the transient regime. Although this analysis needs to be generalized, this result is consistent with astronomical observations of the so-called bullet cluster (with collision speeds on the order of $10^{3}\,\mathrm{km\,s^{-1}}$) and similar objects [32]. These observations put a constraint on the ratio of the collision cross-section to the mass of a purported dark matter (point) particle to $\sigma/M\lesssim 1\,\mathrm{cm^{2}\,g^{-1}}$ [27]. Although the corresponding figure for atomic hydrogen is $\sim 10^{9}\,\mathrm{cm^{2}\,g^{-1}}$, it does not apply in our case. Our analysis does not suggest that atomic hydrogen is a dark matter particle, but rather that the cooperative behavior of a large number of entangled hydrogen atoms is consistent with what is expected for dark matter. While individual atoms may still collide, entangled systems of atoms will be essentially collisionless.
@@ 4 Methods
Because we expect the linear regime response of the system to happen on the superabsorption time-scale $T_{R}$ (see equations 12 and 24) it follows that only a fraction of the atomic population can act cooperatively in the process. That is, given the spectral breadth $\Delta\omega\sim\omega_{0}\Delta v/c$ associated with thermal motions (of characteristic speed $\sim\Delta v$) the density entering our equations is $1\,\mathrm{cm^{-3}}/T_{R}\Delta\omega$. In accordance with the Arecchi-Courtens condition [33], we fixed the thickness of the slab to $L=cT_{R}$, which corresponds to the maximum length of a coherent region when the excitation happens simultaneously over the whole sample (see the corresponding discussion in Sec. 3). Using equation (14) and the Boltzmann equation for the population of energy levels we arrive at $n_{-}^{\mathrm{eq}}=9.6\times 10^{-15}\,\mathrm{cm^{-3}}$, $L=4.5\times 10^{18}\,\mathrm{cm}\simeq 1.5$ pc and $T_{R}=1.5\times 10^{8}\,\mathrm{s}=4.3\times 10^{-7}/\Gamma^{-1}$, with $\Gamma\simeq 2.9\times 10^{-15}\,\mathrm{s^{-1}}$ for the 21 cm line. We therefore set a constant “pump” $\Lambda_{n}=n_{-}^{\mathrm{eq}}/T_{1}$ to enforce thermal equilibrium and keep $n_{-}=n_{-}^{\mathrm{eq}}$ at all times (see equation 52).
@@ Acknowledgements
M.H. conceptualized the work. Both authors contributed equally to the analysis and the writing of the paper.
@@ A.1 Two-atom transition rates from $m=0$ and superabsorption
We consider initial conditions where the states |$\Ket{1,0}_{\theta}$ and |$\Ket{0,0}_{\theta}$ have equal probabilities of occupation (of $1/2$) but with uncorrelated probability amplitude coefficients. Under these conditions, the radiation intensity is given by can be evaluated with
@@
$I\left(t\right)=\frac{1}{2}\hbar\omega_{0}\Gamma e^{-\Gamma t}$$\left\{\left[1+F\left(kz_{0}\right)\right]e^{-\Gamma F\left(kz_{0}\right)t}\right.$
$\left.+\left[1-F\left(kz_{0}\right)\right]e^{\Gamma F\left(kz_{0}\right)t}\right\}.$ (36)

### v4

@@ 2.1 Entangled quantum states and coherent behavior
Following Dicke [1], the kets |$\Ket{r,m}$ are defined with the “cooperative” number $r$ and “inversion” number $m$, which are analogous to quantum numbers for a system of coupled spin-1/2 particles, with $r\geq 0$ and $\left|m\right|\leq r$ (|$\Ket{r,m}_{\theta}=\Ket{r,m}$ when $\cos\theta=0$). Note that $k=2\pi/\lambda$ with $\lambda$ the radiation wavelength.
@@
The two-atom system will, on average, emit a photon half of the time but at twice the single-atom rate. Another way to look at this is by considering the energy trapping efficiency defined as
@@
The calculations can be extended to a larger number of atoms. The general case for an arbitrary number of atoms $n$ is shown in Figure 1 along with the expected transition rates in the small sample (i.e., $kz_{0}\rightarrow 0$) or coherent limits (taken from Dicke’s original paper [1]). It is generally observed that, starting from a given $m$ value, transitions are allowed only between states sharing the same cooperation number $r$, while rates systematically decrease as one moves rightward to lower $r$ values.
@@ 2.3 Superabsorption, coherence and triggered subradiance
We now study the propagation of an incident radiation field on a slab composed of an atomic gas, similar to what is often done in elementary radiative transfer analysis in astrophysics [23]. This will allow us to compare the outcome of our analysis with well-known results. The MBE are perfectly suited for this problem, being one-dimensional in nature. We will also investigate effects the incident field will have on the radiation emanating from within the gas itself. For a magnetic dipole transition, such as the atomic hydrogen 21 cm line, the MBE allow us to follow the evolution of three quantities as a function of position $z$ and the retarded time $\tau=t-z/c$: the differential density $n_{-}=n_{b}-n_{a}$, with $n_{a}$ and $n_{b}$ are the number densities of atoms in the lower and upper levels [^2] From now on we use lower $n$’s for densities and capital $N$’s for numbers of particles., respectively, the magnetization $M^{+}$ and the magnetic field $B^{+}$. Our MBE include relaxation and dephasing time-scales $T_{1}$ and $T_{2}$, as well as a thermal equilibrium condition enforced through a corresponding differential density $n_{-}^{\mathrm{eq}}=n_{b}^{\mathrm{eq}}-n_{a}^{\mathrm{eq}}<0$ set by the Boltzmann equation for the population of energy levels at the temperature of the gas $T_{\mathrm{kin}}$.
@@ 2.3.1 Response to incident radiation
As we will soon be interested in the duration of the transient response of the system, we note that the evolution time-scale of the system is (see Appendix C.2)
$T_{\mathrm{tr}}\approx T_{2}\left(1+\frac{zT_{2}}{LT_{R}}\right)^{-1}$ (15)
when $z\tau\ll LT_{R}$. The transient signal thus dies out on a time-scale of the order of $T_{\mathrm{tr}}$.
@@
$\alpha\approx\alpha_{\mathrm{B}}\left(1+\frac{zT_{2}}{LT_{R}}\right).$ (22)
We uncover the important result that Beer’s law holds strictly only under conditions where $T_{2}\rightarrow 0$, i.e., in the limit of complete decoherence/dephasing. In the opposite limit when $T_{2}\gg T_{R}$, for $z=L$ we have $\delta\omega\sim T_{R}^{-1}$, $\alpha=\alpha_{\mathrm{B}}T_{2}/T_{R}$ and find, just as importantly, that the density of the gas determined with Beer’s law can be overestimated when coherence is present.
@@
Figure 6: Left: Temporal evolutions of the radiation intensity (top) and magnetic induction field (bottom) of the 21 cm line for an atomic hydrogen gas at a temperature of $10$ K and length $L\simeq 1.7$ pc. We set $T_{1}=T_{2}\simeq 14\,T_{R}$ and the intensity is normalized to the saturation intensity $I_{\mathrm{sat}}$ given in equation (18). In the bottom panel, the incident magnetic field’s amplitude is shown in cyan (using the vertical scale on the right), while a broken curve in red corresponding to equation (12) is plotted on top of that obtained from numerical computations (in black) using the MBE. Right: Temporal evolution of the density $n_{-}$ in black with the pump $\Lambda_{n}=n_{-}^{\mathrm{eq}}/T_{1}$ in cyan (using the vertical scale on the right; top) and the magnetization (bottom). The density is limited to the spectral extent of the signal (i.e., $T_{R}^{-1}$) and corresponds to a total density of $1\,\mathrm{cm^{-3}}$ for a gas at a temperature of $10$ K. The mean distance between neighboring cooperating atoms is $\Delta r\approx 740\lambda$ and $T_{R}=1.7\times 10^{8}\,\mathrm{s}\simeq 5.0\times 10^{-7}/\Gamma$.
@@
The left panel of Figure 6 shows the temporal evolutions of the radiation intensity (normalized to $I_{\mathrm{sat}}$; top) and the magnetic field (bottom). A broken curve in red corresponding to equation (12) is plotted on top of that obtained from numerical computations using the MBE for the magnetic field (in black); there is perfect agreement between the two. The (constant) density $n_{-}$ (in black) and pump $\Lambda_{n}$ (in cyan; using the vertical scale on the right; top) and the magnetization (in units of Bohr magneton $\mu_{\mathrm{B}}$ per cubic centimeter; bottom) are displayed in the right panel. We note that the mean distance between neighboring cooperating atoms is $\Delta r\approx 740\lambda$ and $T_{2}\simeq 14T_{R}$, which implies that coherence can be sustained in the system under these conditions.
It is interesting to note that at $\tau=0$ the incident magnetic field is transmitted unattenuated at $z=L$ (or any other position) no matter how large $L$ is. Thereafter, superabsorption defines the transient response to greatly reduce the intensity on a time-scale $\approx T_{R}$ and in a oscillatory manner reminiscing of superradiance systems. According to our previous discussion, the steady-state regime settles afterwards to an intensity diminished by a factor $e^{-\alpha L}$, where $\alpha L=2T_{2}/T_{R}\simeq 28$ from equation (13). Such signal would therefore be detected as a deep and narrow absorption line ($\approx 0.3\,\mathrm{km\,s^{-1}}$ in width) against the background radiation.
@@ 2.3.2 Response to initial fluctuations
More precisely, at $\tau=0$ the system’s intensity is at a level similar to what is expected from a non-coherent gas (i.e., $I\simeq 0.25\,I_{\mathrm{nc}}$), but it only radiates for a duration $\approx T_{R}$. To that effect, we should emphasize that $\Gamma T_{R}\simeq 5.0\times 10^{-7}$. In comparison, for the 21 cm line the non-coherent transient intensity for non-cooperating atoms would last on the order of $\Gamma^{-1}\sim 10^{7}$ yr, whereas $T_{R}\sim 6$ yr. This implies that almost the totality of the available energy is locked within the atomic population, which is almost entirely in dark states since the fast states are quickly depleted during the transient regime.
@@ 3 Discussion
In our analysis of the 21 cm line in Sec. 2.3, we purposely chose a gas density that is in line with measurements in dark matter halos. For the matter contained in these environments, the main source of anisotropy comes in the form of the radiation field emanating from the host galaxy. As then mentioned, the incidence of this field will effectively select one mode for the propagation of radiation within the halo. It is expected that superabsorption of the incident field will elicit coherence in subsequent radiative transitions in the gas through angular correlation among photons. As a consequence, extreme subradiance, leading to a complete darkening of the gas, takes place within a few years even at a mean separation $\Delta r\approx 740\lambda$ between closest neighbors. As far as emission properties are concerned, this atomic hydrogen gas is invisible and effectively behaves like dark matter. This darkening effect may therefore possibly account for the dark matter known to exist in these environments.
@@
For our analysis of the 21 cm line for a gas at a density of $1\,\mathrm{cm^{-3}}$ and temperature $T_{\mathrm{kin}}=10$ K these considerations have no bearing on the outcome since all other lines (e.g., Lyman-$\alpha$) will remain essentially unexcited. The gas as a whole will thus appear dark in the steady-state on the account of the 21 cm line alone. While we only studied this specific example, similar conclusions may be reached when considering other environments with differing gas densities and temperatures, or other atomic/molecular species. It also follows that this darkening effect, which also relies on the presence of anisotropic radiation, should be closely correlated with the presence of “luminous” matter, as is the case for dark matter [27].
@@
We are then led to a picture where our “dark” hydrogen gas (in emission) would also appear more opaque (or more optically thick) in its absorption lines against the background source responsible for the incident radiation. However, it would otherwise be completely transparent to radiation at wavelengths alien to the hydrogen atom, as expected from dark matter. For a cold gas as the one considered in our analysis, these lines would appear as narrow absorption features, such as those commonly detected in observations of the interstellar medium [28], the Galactic halo [29], the Galactic disk [30] and external galaxies [31]. Furthermore, any other field (of weaker intensity) incident on the gas at radiation modes other than the dominant mode responsible for the presence of coherence will propagate unaffected through the gas.
Another important property of our darkened atomic hydrogen gas is discussed in Appendix D. There, we calculate the cross-section for high-speed collisions between two two-atom systems (i.e., when the kinetic energy is far greater than the transition excitation energy $\hbar\omega_{0}$), both being entangled as discussed in Sec. 2.1. Interestingly, it is found that in such circumstances the collision cross-section will likely vanish because of destructive interference due to the entangled nature of the states the systems find themselves in after the transient regime. Although this analysis needs to be generalized, this result is consistent with astronomical observations of the so-called bullet cluster (with collision speeds on the order of $10^{3}\,\mathrm{km\,s^{-1}}$) and similar objects [32]. These observations put a constraint on the ratio of the collision cross-section to the mass of a purported dark matter (point) particle to $\sigma/M\lesssim 1\,\mathrm{cm^{2}\,g^{-1}}$ [27]. Although the corresponding figure for atomic hydrogen is $\sim 10^{9}\,\mathrm{cm^{2}\,g^{-1}}$, it does not apply in our case. Our analysis does not suggest that atomic hydrogen is a dark matter particle, but rather that the cooperative behavior of a large number of entangled hydrogen atoms is consistent with what is expected for dark matter. While individual atoms may collide, entangled systems of atoms will be essentially collisionless.
@@ 4 Methods
Because we expect the linear regime response of the system to happen on the superabsorption time-scale $T_{R}$ (see equations 12 and 24) it follows that only a fraction of the atomic population can act cooperatively in the process. That is, given the spectral breadth $\Delta\omega\sim\omega_{0}\Delta v/c$ associated with thermal motions (of characteristic speed $\sim\Delta v$) the density entering our equations is $1\,\mathrm{cm^{-3}}/T_{R}\Delta\omega$. In accordance with the Arecchi-Courtens condition [33], we fixed the thickness of the slab to $L=cT_{R}$, which corresponds to the maximum length of a coherent region when the excitation happens simultaneously over the whole sample (see the corresponding discussion in Sec. 3). Accounting for the degeneracy of the upper level, using equation (14) and the Boltzmann equation for the population of energy levels we arrive at $n_{-}^{\mathrm{eq}}=7.2\times 10^{-15}\,\mathrm{cm^{-3}}$, $L=5.2\times 10^{18}\,\mathrm{cm}\simeq 1.7$ pc and $T_{R}=1.7\times 10^{8}\,\mathrm{s}=5.0\times 10^{-7}/\Gamma^{-1}$, with $\Gamma\simeq 2.9\times 10^{-15}\,\mathrm{s^{-1}}$ for the 21 cm line. We therefore set a constant “pump” $\Lambda_{n}=n_{-}^{\mathrm{eq}}/T_{1}$ to enforce thermal equilibrium and keep $n_{-}=n_{-}^{\mathrm{eq}}$ at all times (see equation 52).
@@ Acknowledgements
Both authors contributed equally to the analysis and the writing of the paper.
@@
Independence on internal states for high energy scattering D.1
@@ A.1 Two-atom transition rates from $m=0$ and superabsorption
We consider initial conditions where the states |$\Ket{1,0}_{\theta}$ and |$\Ket{0,0}_{\theta}$ have equal probabilities of occupation (of $1/2$) but with uncorrelated probability amplitude coefficients. Under these conditions, the radiation intensity can be evaluated with
@@
$I\left(t\right)=\frac{1}{2}\hbar\omega_{0}\Gamma e^{-\Gamma t}$$\left\{\left[1+F\left(kz_{0}\right)\right]e^{-\Gamma F\left(kz_{0}\right)t}+\left[1-F\left(kz_{0}\right)\right]e^{\Gamma F\left(kz_{0}\right)t}\right\}.$ (36)